give the coordinates of the point where the circle X^2 +y^2=25 intersects the curve 2x^2-3y^2=5
Answer by mananth(16946)

X^2 +y^2=25 intersects the curve 2x^2-3y^2=5
X^2 +y^2=25.......(1)
2x^2-3y^2=5........(2)
multiply eq. (1) by 3 and add to (2)
3x^2+3y^2=75
2x^2-3y^2=5
5x^2=80
x^2=16
x= +/- 4
Plug x=+/- 4 in eq(1)
y^2+16=25
y^2=9
y=+/-3
The points of intersection are
(4,3),(4,-3),(-4,3)(-4,-3)